OpenStudy (anonymous):

if i factor something out of something squared, does the thing i factor out become squared?

OpenStudy (anonymous):

example: ( x + x.5 + x)^2 = x(1 + 1.5 + 1)

OpenStudy (anonymous):

wrong, wait

OpenStudy (anonymous):

\[\sqrt{4x - 4y} = 2\sqrt{x-y}\]

OpenStudy (anonymous):

( x + x.5 + x)^2 = x(1 + 1.5 + 1)^2

OpenStudy (anonymous):

could you use my example?

OpenStudy (anonymous):

Ohhhh. I don't think mine helped.

OpenStudy (anonymous):

so my question is: ( x + x.5 + x)^2 = x(1 + 1.5 + 1)^2 or = x^2(1 + 1.5 + 1)^2

OpenStudy (anonymous):

i'm 99% sure it's the first one

OpenStudy (anonymous):

Is that 0.5x? like x/2?

OpenStudy (anonymous):

i'll write it again for you

OpenStudy (anonymous):

( x + x*5 + x)^2 = x(1 + 1*5 + 1)^2 or = x^2(1 + 1*5 + 1)^2

OpenStudy (anonymous):

im sure it's x(1 + 1*5 + 1)^2, so i'm just double checking

OpenStudy (anonymous):

I think it's x^2(1+5+1)^2. If you were to expand it all terms have x^2 in it.

OpenStudy (anonymous):

Also, (x+5x+x)^2 = (7x)^2

OpenStudy (anonymous):

So, 49x^2

OpenStudy (anonymous):

so it's x^2(1 + 1*5 + 1)^2 ?

OpenStudy (anonymous):

Yeah. The more I think about it, it seems to get more clear. x^2(1+5+1)^2 = (x+5x+x)^2 = 49x^2

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

Yeah. Just try (x+5x+x)(x+5x+x), it's not hard to expand out. Thats what I did to make sure I was telling you the right thing :)

OpenStudy (anonymous):

alright, thanks for clearing that up